OpenStudy (anonymous):

2x^4=10 Would I square root it twice? Or?

OpenStudy (anonymous):

\[2x^{4} = 10\] x * x * x *x = x^4

OpenStudy (anonymous):

I understand that, not sure how that gets me anywhere though,

OpenStudy (anonymous):

well we can divide both sides by 2

OpenStudy (anonymous):

So x^4 = 5

OpenStudy (anonymous):

yeah then \[\sqrt[4]{5} = x\]

OpenStudy (anonymous):

Is that the answer? Or is it an decimal?